what is the graph answer to -x^2 + -14x = 49
Answer by ewatrrr(24785)


Hi
Graphing:
-x^2 + -14x = 49
-(x^2 +14) = 49
-[(x+7)^2 -49] = 49
-(x+7)^2 + 49 = 49
-(x+7)^2 = 0
(x+7)^2 = 0
f(x) = (x+7)^2 Vertex is Pt(0,-7)